If the terminal side of a 330 degree angle intersects a unit circle, what would the coordinates at the point of intersection?
Answer by Alan3354(69443)

If the terminal side of a 330 degree angle intersects a unit circle, what would the coordinates at the point of intersection?
--------
x = cos(330)
y = sin(330)
==================
For any angle A:
x = cos(A)
Y = sin(A)